Joanne is 18 and her mother is 43. In how many years will her mother be twice as old as Joanne ?

Answer by JulietG(1812)

There are many ways to do this problem. Here's one way:
Twice Joanne's age is 36. Subtract that from 43.
43 - 36 = 7
In 7 years, the mother will be twice as old.
Let's check.
2(18 + 7) = 43+7
2(25) = 50
50=50